OpenStudy (anonymous):

How do I solve the systems of equations by substitution 5x-16=-y,3x-y=0

OpenStudy (anonymous):

5x - 16 = -y 3x - y = 0 solve for y for either one above then substitute into the other equation 5x -16 = -y y = 16 - 5x substitute this into 3x - y = 0

OpenStudy (anonymous):

Thanks Gaby and nphoungsun 93.